# Applications of Spiking Neural Networks in Visual Place Recognition

### v2

@@ IV-C Proof-of-concept Robot Deployment Setup
For our proof-of-concept CPU-based robot deployment experiment of our Modular SNN, we used AgileX’s Scout Mini robot [60] to navigate the QUT Centre for Robotics floor. This area is a shared space with other researchers, where there are moving people, and robots, as well as relocated objects. The robot was tele-operated at a speed of approximately 1 m/s. It collected reference and query traverses at a frequency of 1 Hz at different times of the day, encountering challenges such as occlusions, brightness changes, and slight lateral and frontal viewpoint changes. In this experiment, we used a 32GB Intel Core i7 CPU for processing, and an image-based camera (Intel’s RealSense D435) for capturing images of the environment.
@@ IV-E Baseline Methods
We employ several conventional VPR approaches to evaluate the performance of our methods. These techniques encompass a non-learned method: Sum-of-Absolute-Differences (SAD) [41], as well as learning-based methods: DenseVLAD [48], NetVLAD [49], AP-GeM [50], Generalized Contrastive Loss (GCL) [51], CosPlace [52], and MixVPR [53]. We also include a comparison to a previous Non-modular SNN [59] for VPR. This comparison is included to show the scalability advantages of ensembling and modularity in terms of both performance and computational time as the number of places to learn increases. These approaches are detailed as follows:
@@
Non-modular SNN [59]: is a single three-layer SNN model that is enlarged to accommodate for learning a significantly larger number of places. The model is trained on the reference dataset, and tested on the query dataset.
@@ V-A Component-Wise Contributions to SNN Performance
This section overviews the performance contributions of each component of our approach, namely modularity (i.e. expert modules that learn small subsets of the reference dataset), ensembling (i.e. representing each place by multiple modules), and sequence matching (i.e. using multiple reference and query images for place matching). Figure 4 demonstrates that the performance of our Modular SNN is significantly increased when each of these techniques, ensembling (R@1 increase of 6.3%) and sequence matching (R@1 increase of 17.2% for a sequence length of four), is applied separately. Moreover, the combination of ensembling and sequence matching techniques further improves the R@1 of the Modular SNN (R@1 increase of 24.9%). As the ensembling and sequence matching techniques are commutative, the order of application of these two techniques produces identical outcomes. The Non-modular SNN by Hussaini et al. [59] performs poorly, even when ensembling and sequence matching are applied (R@1 is less than 0.2%).
@@ V-B Comparison of Ensemble of Modular SNNs without a Sequence Matcher to Conventional VPR Techniques
Figure 5: R@1 performance improvements with sequence matching: The plot shows the mean R@1 performance of each method across all datasets when employing sequence matching using four, seven and ten frames, compared to the single-frame approach (SL1). The gray lines represent the standard deviation of the R@1 model performance across all datasets. The red bar demonstrates the mean R@1 performance improvement of a method without a sequence matcher to the performance of the method with a sequence matcher of sequence length ten. Our Modular and Ensemble of Modular SNNs obtain the highest R@1 improvement with a sequence matcher (from without a sequence matcher to a sequence matcher of sequence length ten). The mean R@1 performance of both our Ensemble of Modular SNNs, and Modular SNN without a sequence matcher (SL1) is competitive with multiple VPR techniques, and incorporating a sequence matcher with sequence lengths of four, seven, and ten enables our SNN-based approaches to obtain the highest R@1 improvement compared to similar-performing VPR baselines. Notably, the R@1 performance of our SNN-based approaches with a sequence matcher of sequence length ten frames slightly surpasses that of AP-GeM (ResNet101, LM18), and is in close approximately to that of DenseVLAD, both of which have higher-performing baselines (without a sequence matcher).
@@ V-C Comparison of Ensemble of Modular SNNs with a Sequence Matcher to Conventional VPR Techniques
Compared to VPR techniques with roughly similar baseline performance, our Ensemble of Modular SNNs obtains the overall highest improvement with a sequence matcher averaged across all datasets: the mean R@1 of our Ensemble of Modular SNNs with a sequence matcher across all datasets is among the top five models, despite the lower baseline performance without a sequence matcher. Specifically, Our Ensemble of Modular SNNs with a four-frame sequence matcher achieved a 39.1% R@1 performance gain on the SFU Mountain dataset with a baseline performance of 46.1%, surpassing the 22.9% increase of the comparable VPR technique, GCL, with a similar baseline performance of 47.2%.
@@
Figure 6: Indication of sequence matching responsiveness: The figure shows the correct match sparsity against the R@1 performance ratio of methods with a sequence matcher of sequence lengths four to one. Correct match sparsity is defined as the mean distance to the next correct match for all predictions of query images, projected to log space and min-max normalized for each dataset to allow for dataset agnostic evaluation. Higher sparsity indicates wider gaps between correct matches, suggesting potential under-performance. The gray line represents the best fit line for all data points. Key observations are: 1) Our Modular SNN and Ensemble of Modular SNNs generally show a higher SL4 to SL1 ratio compared to nearby data points, regardless of the dataset the methods were evaluated on, indicating strong adaptability to sequence matching. Exceptions include our Ensemble’s performance on Nordland FS, Synthia Night to Fall, and our Modular SNN evaluated on Nordland FS, which are comparable to other methods. 2) Across all data points, methods with high correct match sparsity show greater responsiveness to sequence matching, regardless of the method observed.
@@ V-H Proof-of-Concept Robot Deployment
Figure 9 shows the reference path taken by the robot, with correct and incorrect predictions of our Modular SNN at query time marked by green and red markers, respectively. The figure indicates that the model generally performs well, with R@1 = 75.0%. Most misclassifications occurred at curves or where the query image position is slightly shifted laterally and/or frontally. This is due to the limitation in viewpoint tolerance of our approach, which is discussed in Section VI as part of our future work.
We used a ground truth tolerance of 3 m, due to the high visual overlap between consecutive places. The inference time of our Modular SNN ranged from 1.1 to 2 seconds per image. This experiment validated the feasibility of our Modular SNN approach for real-time robot deployment in a small indoor environment.
@@
Figure 9: Robot deployment feasibility study: [Left] Real-time deployment of our Modular SNN in a small indoor environment on a CPU. First, the reference dataset is collected and the Modular SNN is trained offline. During inference, the robot moves through the environment, collecting images, and predicting their place labels in real-time. The reference and query images are collected at different times of the day. The images were collected at 1 Hz with the robot moving at approximately 1 m/s. The blue line and points represent the reference path and images, while the green line represents the query path. Correct predictions are marked with green crosses, and incorrect predictions are marked with red crosses. The four samples of the query and predicted places show the original images and their preprocessed forms, which are used as input to our Modular SNN. [Right] Proof-of-concept robot deployment testing platform. We used an AgileX Scout Mini [60] robot equipped with a Intel’s RealSense D435 camera.
Figure 10: Qualitative results: The plot showcases the performance of our Standalone Modular SNN, Ensemble of Modular SNNs, and various VPR methods across diverse datasets. To enhance clarity, we have included just one of the three NetVLAD instances, specifically NetVLAD (Landmarks), and one of the three AP-GeM instances, namely AP-GeM (ResNet101, LM18). It details three instances of correct predictions by both Modular SNN and its Ensemble variant (in rows one, four, and five), two cases where the Ensemble yields correct matches despite the incorrect prediction of the Modular SNN as an ensemble member (in rows two and six), and a situation where both Modular and Ensemble of Modular SNNs fail to correctly match the query image to its corresponding reference image (in row three).
@@ VI Discussions and Conclusions
Lastly, in addition to ensembling, we also explored the impact of sequence matching, a technique that further augments our system’s performance by utilizing multiple consecutive images for place matching. Pairing our Ensemble of Modular SNNs with sequence matching exhibited a higher R@1 performance improvement compared to VPR techniques with similar baselines, except in the case of the Oxford RobotCar dataset. This reinforces the transformative role of sequence matching on enhancing SNN capabilities for visual place recognition. We also provided an indicator for sequence matching responsiveness, applicable to general VPR techniques, which demonstrated the competitive adaptability of our SNN-based solutions to sequence matching compared to that of VPR techniques.
Our work follows the similar trend seen in the recent state-of-the-art conventional VPR techniques such as [52] that frame the visual place recognition problem as a classification task, enabling large-scale recognition capability by bypassing the computationally heavy process of computing the pairwise distance matrix for all query and reference feature representations. Another benefit of approaching the VPR problem as a classification task is that it eliminates the need to divide the reference and query datasets into separate training and testing partitions, effectively circumventing the potential complication of overlapping data in the training and testing subsets [52]. We highlight that our approach is trained only on the reference set, which has geographical overlap with the query set of the same dataset used for testing. This strategy is advantageous for real-world applications as it avoids pretraining on large datasets and uses fewer training images compared to conventional VPR techniques, which rely on extensive training datasets that are typically geographically separate from the reference and query set of the dataset used for evaluation. Furthermore, we demonstrate the performance of our SNN approach using low resolution image sizes, which scales well with increasing number of places in terms of storage complexity. In comparison, conventional VPR techniques typically need to store reference image feature descriptors, posing challenges for scaling to large datasets due to associated increase in storage requirements.
@@
A limitation of our work is the minimal robustness to viewpoint shift, an aspect that most VPR techniques address effectively. Enhancing the resilience of our system to viewpoint change remains a priority for us, as it is crucial for reliable place recognition in more challenging situations. To overcome this challenge, we can incorporate an attention-based mechanism similar to recent SOTA Transformer-based VPR approaches [149, 150], or divide each input image into smaller patches, each processed by a module, similar to regional descriptors such as PatchNetVLAD [128]. However, the necessity for robustness against significant viewpoint shifts may vary depending on the specific downstream application of our visual place recognition system. For instance, in scenarios where the VPR system acts as a loop closure component of a Simultaneous Localization and Mapping (SLAM) process, the limitations of the SLAM system in loop closure might render extreme viewpoint robustness less critical [20].

### v3

@@ IV-C Proof-of-concept Robot Deployment Setup
For our proof-of-concept CPU-based robot deployment experiment of our Modular SNN, we used AgileX’s Scout Mini robot [60] to navigate the QUT Centre for Robotics floor. This area is a shared space with other researchers, where there are moving people, and robots, as well as relocated objects. The robot was tele-operated at a speed of approximately 1 m/s. It collected reference and query traverses at a frequency of 1 Hz at different times of the day, encountering challenges such as occlusions, brightness changes, and slight lateral and frontal viewpoint changes. In this experiment, we used a 32GB Intel Core i7 CPU for processing, and Intel’s RealSense D435 for capturing images of the environment.
@@ IV-E Baseline Methods
We employ several conventional VPR approaches [41, 48, 49, 50, 51, 52, 53] to evaluate the performance of our methods, as well as a comparison to a previous Non-modular SNN [59] for VPR. These approaches are detailed as follows:
@@
Non-modular SNN [59]: is a single three-layer SNN model that is enlarged to accommodate for learning a significantly larger number of places. The model is trained on the reference dataset, and tested on the query dataset. This comparison is included to show the scalability advantages of ensembling and modularity in terms of both performance and computational time as the number of places to learn increases.
@@ V-A Component-Wise Contributions to SNN Performance
This section overviews the performance contributions of each component of our approach, namely modularity (i.e. expert modules that learn small subsets of the reference dataset), ensembling (i.e. representing each place by multiple modules), and sequence matching (i.e. using multiple reference and query images for place matching). Figure 4 demonstrates that the performance of our Modular SNN is significantly increased when each of these techniques, ensembling (R@1 increase of 6.3% with five ensemble members) and sequence matching (R@1 increase of 17.2% for a sequence length of four), is applied separately. Moreover, the combination of ensembling and sequence matching techniques further improves the R@1 of the Modular SNN (R@1 increase of 24.9%). As the ensembling and sequence matching techniques are commutative, the order of application of these two techniques produces identical outcomes. The Non-modular SNN by Hussaini et al. [59] performs poorly, even when ensembling and sequence matching are applied (R@1 is less than 0.2%).
@@ V-B Comparison of Ensemble of Modular SNNs without a Sequence Matcher to Conventional VPR Techniques
Figure 5: R@1 performance improvements with sequence matching: The plot shows the mean R@1 performance of each method across all datasets when employing sequence matching using four, seven and ten frames, compared to the single-frame approach (SL1). The gray lines represent the standard deviation of the R@1 model performance across all datasets. Red bars demonstrate the mean R@1 performance improvement of a method without a sequence matcher to the performance of the method with a sequence matcher of sequence length ten. Our Modular and Ensemble of Modular SNNs obtain the highest R@1 improvement with a sequence matcher (from without a sequence matcher to a sequence matcher of sequence length ten). The mean R@1 performance of both our Ensemble of Modular SNNs (with five ensemble members), and Modular SNN without a sequence matcher (SL1) is competitive with multiple VPR techniques, and incorporating a sequence matcher with sequence lengths of four, seven, and ten enables our SNN-based approaches to obtain the highest R@1 improvement compared to similar-performing VPR baselines. Notably, the R@1 performance of our SNN-based approaches with a sequence matcher of sequence length ten frames slightly surpasses that of AP-GeM (ResNet101, LM18), and is in close approximately to that of DenseVLAD, both of which have higher-performing baselines (without a sequence matcher).
@@ V-C Comparison of Ensemble of Modular SNNs with a Sequence Matcher to Conventional VPR Techniques
Compared to VPR techniques with roughly similar baseline performance, our Ensemble of Modular SNNs (with five ensemble members) obtains the overall highest improvement with a sequence matcher averaged across all datasets: the mean R@1 of our Ensemble of Modular SNNs with a sequence matcher across all datasets is among the top five models, despite the lower baseline performance without a sequence matcher. Specifically, our Ensemble of Modular SNNs with a four-frame sequence matcher achieved a 39.1% R@1 performance gain on the SFU Mountain dataset with a baseline performance of 46.1%, surpassing the 22.9% increase of the comparable VPR technique, GCL, with a similar baseline performance of 47.2%.
@@
Figure 6: Indication of sequence matching responsiveness: The figure shows the correct match sparsity against the R@1 performance ratio of methods with a sequence matcher of sequence lengths four to one. Correct match sparsity is defined as the mean distance to the next correct match for all predictions of query images, projected to log space and min-max normalized for each dataset to allow for dataset agnostic evaluation. Higher sparsity indicates wider gaps between correct matches, suggesting potential under-performance. The gray line represents the best fit line for all data points. Key observations are: 1) Our Modular SNN and Ensemble of Modular SNNs (with five ensemble members) generally show a higher SL4 to SL1 ratio compared to nearby data points, regardless of the dataset the methods were evaluated on, indicating strong adaptability to sequence matching. Exceptions include our Ensemble’s performance on Nordland FS, Synthia Night to Fall, and our Modular SNN evaluated on Nordland FS, which are comparable to other methods. 2) Across all data points, methods with high correct match sparsity show greater responsiveness to sequence matching, regardless of the method observed.
@@ V-H Proof-of-Concept Robot Deployment
Figure 9 shows the reference path taken by the robot, with correct and incorrect predictions of our Modular SNN at query time marked by green and red markers, respectively. The figure indicates that the model generally performs well, with a R@1 of 75.0%. Most misclassifications occurred at curves or where the query image position is slightly shifted laterally and/or frontally. This is due to the limitation in viewpoint tolerance of our approach, which is discussed in Section VI as part of our future work.
We used a ground truth tolerance of 3m, due to the high visual overlap between consecutive places. The inference time of our Modular SNN ranged from 1.1 to 2.0 seconds per image. This experiment validated the feasibility of our Modular SNN approach for real-time robot deployment in a small indoor environment.
@@
Figure 9: Robot deployment feasibility study: Left: Real-time deployment of our Modular SNN in a small indoor environment on a CPU. First, the reference dataset is collected and the Modular SNN is trained offline. During inference, the robot moves through the environment, collecting images, and predicting their place labels in real-time. The reference and query images are collected at different times of the day. The images were collected at 1 Hz with the robot moving at approximately 1 m/s. The blue line and points represent the reference path and images, while the green line represents the query path. Correct predictions are marked with green crosses, and incorrect predictions are marked with red crosses. The four samples of the query and predicted places show the original images and their preprocessed forms, which are used as input to our Modular SNN. Right: Proof-of-concept robot deployment testing platform. We used an AgileX Scout Mini [60] robot equipped with an Intel RealSense D435 camera.
Figure 10: Qualitative results: The plot showcases the performance of our Standalone Modular SNN, Ensemble of Modular SNNs (with five ensemble members), and various VPR methods across diverse datasets. To enhance clarity, we have included just one of the three NetVLAD instances, specifically NetVLAD (Landmarks), and one of the three AP-GeM instances, namely AP-GeM (ResNet101, LM18). It details three instances of correct predictions by both Modular SNN and its Ensemble variant (in rows one, four, and five), two cases where the Ensemble yields correct matches despite the incorrect prediction of the Modular SNN as an ensemble member (in rows two and six), and a situation where both Modular and Ensemble of Modular SNNs fail to correctly match the query image to its corresponding reference image (in row three).
@@ VI Discussions and Conclusions
Lastly, in addition to ensembling, we also explored the impact of sequence matching, a technique that further augments our system’s performance by using multiple consecutive images for place matching. Pairing our Ensemble of Modular SNNs with sequence matching exhibited a higher R@1 performance improvement compared to VPR techniques with similar baselines, except in the case of the Oxford RobotCar dataset. This reinforces the significant role of sequence matching on enhancing SNN capabilities for visual place recognition. We also provided an indicator for sequence matching responsiveness, applicable to general VPR techniques, which demonstrated the competitive adaptability of our SNN-based solutions to sequence matching compared to that of VPR techniques.
Our work follows the similar trend seen in the recent state-of-the-art conventional VPR techniques such as [52] that frame the visual place recognition problem as a classification task, enabling large-scale recognition capability by bypassing the computationally heavy process of computing the pairwise distance matrix for all query and reference feature representations. We highlight that our approach is trained only on the reference set, which has geographical overlap with the query set of the same dataset used for testing. This strategy is advantageous for real-world applications as it avoids pretraining on large datasets and uses fewer training images compared to conventional VPR techniques, which rely on extensive training datasets that are typically geographically separate from the reference and query set of the dataset used for evaluation. Furthermore, we demonstrate the performance of our SNN approach using low resolution image sizes, which scales well with increasing number of places in terms of storage complexity. In comparison, conventional VPR techniques typically need to store reference image feature descriptors, posing challenges for scaling to large datasets due to associated increase in storage requirements.
@@
A limitation of our work is the minimal robustness to viewpoint shift, an aspect that most VPR techniques address effectively. Enhancing the resilience of our system to viewpoint change remains a priority for us, as it is crucial for reliable place recognition in more challenging situations. To overcome this challenge, we can incorporate an attention-based mechanism similar to recent transformer-based VPR approaches [149, 150], or divide each input image into smaller patches, each processed by a module, similar to regional descriptors such as Patch-NetVLAD [128]. However, the necessity for robustness against significant viewpoint shifts may vary depending on the specific downstream application of our visual place recognition system. For instance, in scenarios where the VPR system acts as a loop closure component of a Simultaneous Localization and Mapping (SLAM) process, the limitations of the SLAM system in loop closure might render extreme viewpoint robustness less critical [20].